6. I have created a triangular garden such that the largest side is 8m less than twice the smallest
and the medium side is 12m larger than the smallest side. If the total perimeter of the garden is
104m, what are the lengths of the three sides?

Answer by addingup(3677)

2x-8+x+12+x = 104
4x+4 = 104
4x = 100
x = 25 is the small side.
The medium side: 25+12 = 37
The largest side: 2(25)-8 = 50-8= 42
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
25+37+42 = 104 we have the correct answer.
J